We have got basal material – University of Copenhagen

Forward this page to a friend Resize Print Bookmark and Share

NEEM > Field diaries > Field season 2012 > We have got basal mate...

08 June 2012

We have got basal material

An example of the basal material retrieved today. All the tiny dark spots are mineral particles. Using luminescence technique, we may determine the time when they last were exposed to sunlight.

 

One day after her arrival, Christine is already happy to have obtained a series of samples central to her research. The drillers were able to retrieve basal material from 2538 m depth and it was extracted under ideal conditions for Christine’s work with luminescence dating. A fine success.

Although our weather forecast hinted reasonable conditions, it was not possible today to launch a drilling mission with the Basler (DC-3). It has been cloudy and snowy all day. The group scheduled for drilling instead continued excavation of the mechanic garage.  

A group worked on laying roof over the entrance to the experimental trench, later they began to dig out all the excess snow that had been accumulated during the burial of the balloon. As the end of the balloon was exposed, it turned out that the snow had packed and hardened nicely in the last three days. It was decided to deflate the balloon, and as the balloon deflated, a large cavern of snow was exposed. Most of the floor is the original flat bottom of the trench; but the ceiling is vaulted as it was cast around the cylindrical balloon.

What we have done today:
1. Drilled with HT drill with hardened cutters in deep hole.
2. Today’s Basler mission cancelled.
3. NEGIS team preparing seismic profile West of NEEM camp.
4. Grooming skiway with beam groomer and resetting skiway markers.
5. Preparing drill trench for extraction of basal material in red light.
6. Laying roof on the inclined entrance to the experimental trench.
7. Deflating balloon in experimental trench.
8. Excavation of the mechanic garage almost complete.
9. Snowmobile with lost ski repaired.

Ad.1 and Ad.4: Drillers made drilling efforts with hardened cutters and did not have good penetration. Before the drill came to the surface, all lights were switched off and dark room lighting was established. No core was retrieved; but the chip chamber contained substantial amounts of chips with basal particles. The chips were retrieved and packed in light tight packing in dark room light. The particles in the chips will later be analyzed with stimulated luminescence for dating the basal layers. When the light was switched on again, drillers could see that the cutters were damaged. They have decided to switch to the rock drill and will attempt drilling with this tomorrow.

Ad.6 : This evening the balloon the in the experimental trench was deflated, and access was gained. The roof over the trench is pure compacted snow, and it is holding out fine. The balloon has made a smooth vaulted roof and the snow cast around the balloon is very hard. Tomorrow the new trench will be inaugurated and measurements of roof height, width etc. will be recorded. It is our hope that it will be possible to measure how the trench compresses over time in the coming years.

Weather: Overcast with snow showers all day. Temp. – 7°C to – 16°C, 4 - 14 knots from S. Visibility: 1 km during snow showers, otherwise unrestricted.

FL, J.P. Steffensen



← Previous entry                                                                              Next entry →